OpenStudy (anonymous):

A line has a slope of −5 and contains the point (−2, 4). Which equations represent the line? Choose all answers that are correct. A. −5x + y = −6 B. 5x + y = −6 C. (y – 4) = −5(x + 2) D. (y + 4) = −5(x − 2)

OpenStudy (anonymous):

it is d

OpenStudy (anonymous):

because the formula for point slope is y-y1=slope(x-x1)

OpenStudy (chunkymonkay):

its only d @zynsaarmour

OpenStudy (chunkymonkay):

@zynsarmour

OpenStudy (chunkymonkay):

@StudyGurl14

OpenStudy (studygurl14):

First put in point-slope to get the first equation. then take that and convert it to slope-intercept form to find the second equation. Point slope: \(\Large (y-y_1)=m(x-x_1)\) with point \((x_1,y_1)\) on the line, and m=slope.

OpenStudy (chunkymonkay):

thanks